John is four times as old as Harry. In six years, John will be twice as old as Harry. What are their ages now?
Answer by Earlsdon(6103)

Let J = John's present age and H = Harry's present age.
From the problem description, you can write...
1) J = 4H "John is four times as old as Harry" and...
2) J+6 = 2(H+6) "In six years, John will be twice as old as Harry (will be in six years)"
Substitute the J = 4H from equation 1) into equation 2) and solve for H.
2a) (4h)+6 = 2(H+6) Simplify.
2a) 4H+6 = 2H+12 Subtract 2H from both sides.
2a) 2H+6 = 12 Subtract 6 from both sides.
2a) 2H = 6 Finally, divide both sides by 2.
2a) H = 3 and...
1) J = 4(3) = 12
John is 12 years old and Harry is 3 years old.